What is an equation in standard form with integer coefficients for the line with a slope 8/7 through point (-4,-5)?
Answer by algebrahouse.com(1659)

Point-slope form is y - y1 = m(x - x1)
m is the slope
(x1 , y1) are the coordinates of the point

(-4 , -5) , slope = 8/7
y - (-5) = (8/7)[x - (-4)] {substituted point and slope into point-slope form}
y + 5 = (8/7)(x + 4) {simplified double negative signs}
7y + 35 = 8(x + 4) {multiplied entire equation by 7 to eliminate fraction}
7y + 35 = 8x + 32 {used distributive property}
8x - 7y = 3 {subtracted 7y and 32 from each side to get into standard form}

For more help from me, visit: www.algebrahouse.com